OpenStudy (anonymous):

Simplify 5√112 A) 80√7 B) 9√7 C) 20√7 D) 4√7

OpenStudy (anonymous):

\[112=16\times 7\] so \[\sqrt{112}=\sqrt{16\times 7}=\sqrt{16}\times \sqrt{7}=4\times \sqrt{7}\]

OpenStudy (anonymous):

therefore you get \[20\sqrt{7}\]

OpenStudy (anonymous):

i thank C